A 40.0-kg child running at a speed 3.00 m/s jumps on a stationary playground merry-go-round at a distance 1.50 m from the axis of rotation of the merry-go-round. The child is traveling tangential to the edge of the merry-go-round which has a 600 kg.
Answer #1

Rotational momentum of child = I?

I = mr2 = 40 x 1.52 = 90 kg-m2

tangential velocity = 3 m/s = ? x r

3/r = 3/1.5 = ? = 2 rad/s

momentum of child = 90 x 2 = 180 kg-m2/s

Momentum before = momentum after

180 kg-m2/s = (90 + 600)?

180/(90 + 600) = ? = 0.26087 rad/s